Three squares have sides of different integer lengths, a, b, and c. The total area of the three squares cannot be
(A) 14 (B) 29 (C) 50 (D) 88 (E) 101

Answer by checkley77(12844)

1+4+9=14
4+9+16=29
9+16+25=50
1+36+64=101
ANSWER D)88.
(A) 14 (B) 29 (C) 50 (D) 88 (E) 101